OpenStudy (anonymous):

Find the variation constant and an equation of variation where y varies directly as x and y= 32 when x =8

OpenStudy (saifoo.khan):

y = kx

OpenStudy (saifoo.khan):

where k is a constant.

OpenStudy (saifoo.khan):

Find the value of k. 32 = k* 8 k = 32/8 k = 4 Final equation: y = 4x.

OpenStudy (anonymous):

saifoooooooooooo!

OpenStudy (saifoo.khan):

Sateliteeeeeeeeeee!